without using a calculator or a table, find the value of each expression. leave your answers in terms of pi whenever appropriate.
a. inverse tan of 1
b. inverse sin of sqrt 3/2
c.invers cos (-sqrt 3/2)
d.inverse tan sqrt 3
e.inverse sin of 0
f. sin(inverse cos 3/5)
g. sec(inverse tan 5/12)
h. cos(inverse sin of 1)

Answer by venugopalramana(3286)

without using a calculator or a table, find the value of each expression. leave your answers in terms of pi whenever appropriate.
a. inverse tan of 1...THAT IS RIGHT ISOCELLES TRIANGLE............PI/4
b. inverse sin of sqrt 3/2...THAT IS HALF EQUILATERAL TRIANGLE..PI/3
c.invers cos (-sqrt 3/2).THAT IS SECOND Q.AND HALF EQUILATERAL TRIANGLE.....PI-PI/6=5PI/6
d.inverse tan sqrt 3...THAT IS HALF EQUILATERAL TRIANGLE=PI/3
e.inverse sin of 0..=0
f. sin(inverse cos 3/5)=THAT IS COS(X)=3/5...HENCE SIN(X)=SQRT(1-COS^2(X))=4/5
g. sec(inverse tan 5/12)..THAT IS TAN(X)=5/12...HENCE SEC(X)=SQRT(1+TAN^2(X))=13/12
h. cos(inverse sin of 1)...THAT IS SIN(X)=1...HENCE COS(X)=SQRT(1-SIN^2(X))=0